# Quasi-steady emission from repeating fast radio bursts can be explained by magnetar wind nebula

### v1

@@ 2.2.1 Energy injection
Both rotational energy injection (Dai et al., 2017; Kashiyama & Murase, 2017) and magnetic energy injection (Beloborodov, 2017; Margalit & Metzger, 2018; Zhao & Wang, 2021) models have been proposed to explain the PRS luminosity as well as the large RM detected for FRB 121102 (Michilli et al., 2018). The observed persistent radio emission from FRBs can be interpreted as synchrotron emission from electrons in the nebula.
@@
[…]
@@ 2.2.2 Particle number distribution
The rotational or magnetic energy is injected into the nebula together with the particles. The injected particles are accelerated to relativistic energies inside or around the termination shock before entering the nebula. The lepton injection rate $\dot{n}_{\mathcal{E}_{e}}^{\rm inj}$ is given by a broken power-law function (see e.g, Tanaka & Takahara 2010; Murase et al. 2015)
$\mathcal{E}_{e}^{2}\dot{n}_{\mathcal{E}_{e}}^{\rm inj}=\frac{3\epsilon_{e}(L_{\rm sd}+L_{\rm mag})}{4\pi R_{\rm nb}^{2}c\mathcal{R}_{0}}\left\{\begin{array}[]{ll}(\gamma_{e}/\gamma_{b})^{2-q_{1}},&\gamma_{e}\leq\gamma_{b}\vspace{0.2cm}\\ (\gamma_{e}/\gamma_{b})^{2-q_{2}},&\gamma_{b}<\gamma_{e}\\ \end{array}\right.$ (4)
where $R_{\rm nb}$ is the nebula radius, $\mathcal{R}_{0}\sim(2-q_{1})^{-1}+(q_{2}-2)^{-1}\sim 5$ is the correction factor for the lepton normalization, $q_{1}<2$ and $q_{2}\geq 2$ are the low- and high-energy spectral indices. For our analysis, we adopt the characteristic LF of the accelerated leptons to be $\gamma_{b}\sim 10^{5}$ for rotation-powered and $\gamma_{b}\sim 10^{3}$ for magnetic-powered models. The spectral indices are $q_{1}=1.5$ and $q_{2}=2.5$ for rotation-powered scenario, and $q_{1}=q_{2}=2.0$ for magnetic-powered case. We assume $10^{3}\leq\gamma_{e}\leq 10^{7}$, based on the multi-wavelength modelling of young Galactic PWNe (Tanaka & Takahara, 2013). The effective pair multiplicity can be generally expressed in terms of $\gamma_{b}$, $q_{1}$ and $q_{2}$ (see Murase et al. 2015).
@@ 3.1 MWN and SNR evolution
The density profile of the SN ejecta can be described using a power-law function $\rho_{\rm ej}=(3-\delta)M_{\rm ej}/4\pi R_{\rm ej}^{3}(R/R_{\rm ej})^{-\delta}$ (Chevalier & Soker, 1989; Kasen & Bildsten, 2010), where $\delta=1$ is the fiducial value of the index, $M_{\rm ej}$ is the ejecta mass and $R_{\rm ej}$ is the ejecta radius. The time evolution of the SN ejecta internal energy is given by
@@ 3.2 Synchrotron emission
The average magnetic field in the nebula is given by $B_{\rm nb}=\sqrt{6\mathcal{E}_{B}/R_{\rm nb}^{3}}$, where $\mathcal{E}_{B}$ is the magnetic energy in the nebula. The magnetic energy in the nebula evolves as (see e.g., Murase et al. 2016; Murase et al. 2021)
@@
In this work, we consider $c_{B}=0$ in the limit when energy loss due to adiabatic expansion is negligible, as done previously for Galactic PWNe (Tanaka & Takahara, 2010) and pulsar-powered PWNe (Murase et al., 2015, 2016).
We compute the synchrotron emission from the nebula by solving kinetic equations for the photons and electrons (see e.g., Murase et al. 2021)
@@
where $t_{\rm comp}^{\rm nb}$ is the Comptonisation timescale in the nebula, $t_{\rm esc}^{\rm nb}=R_{\rm nb}/c$ is the photon escape timescale, $t_{\gamma\gamma}$ is the photon annihilation timescale, $\partial n_{\mathcal{E}_{\gamma}}^{\rm IC/syn}/\partial t$ is the photon generation rate from IC/synchrotron process. $P_{\rm ad}$, $P_{\rm syn}$ and $P_{\rm IC}$ are the energy-loss rates due to adiabatic expansion, synchrotron radiation and IC, respectively, $\partial n_{\mathcal{E}_{e}}^{\gamma\gamma}/\partial t$ is the electron production rate from photon annihilation and $\dot{n}_{\mathcal{E}_{e}}^{\rm inj}$ is the electron injection rate from equation (4).
@@ 4.1 Constraints on model parameters
Figure 2: Constraints on NS parameters $B_{\rm dip}$ and $P_{i}$ from the nebula energy requirement (red curves), ejecta and nebula contribution to the source DM (green curves), and the NS spindown luminosity (black curves) are shown for FRB 121102 in top row, FRB 190520 in middle row and FRB 201124 in bottom row. In each panel, the solid curves show the results for ($E_{\rm sn}$, $M_{\rm ej}$) = ($10^{50}\,{\rm erg}$, $0.1\,M_{\odot}$) while the dashed curves show results for ($E_{\rm sn}$, $M_{\rm ej}$) = ($10^{51}\,{\rm erg}$, $3.0\,M_{\odot}$). For the magnetic energy injection, we fix $B_{\rm int}=10^{16}\,{\rm G}$ and $t_{\rm inj}=0.6\,{\rm yr}$ for all cases. For FRBs 121102 and 190520, we vary $t_{\rm age}=10\,{\rm yr}$ (left panel), $20\,{\rm yr}$ (middle panel) and $40\,{\rm yr}$ (right panel) as $t_{\rm obs}>5\,{\rm yr}$. In case of FRB 201124, results are shown for $t_{\rm age}=5\,{\rm yr}$ (left panel), $10\,{\rm yr}$ (middle panel) and $20\,{\rm yr}$ (right panel).
  • Magnetised nebula energy requirement: The minimum energy in relativistic electrons required to explain the quasi-steady radio emission from FRB sources should be less than the maximum energy stored in the nebula which is given by ${\rm Max}(\mathcal{E}_{e}^{\rm inj})$ (see e.g., equation 2.2.1). The observed quasi-steady radio spectrum for FRBs can be fitted by
    $\nu F_{\nu}\approx\frac{\mathcal{L}_{\rm nb}}{4\pi d_{L}^{2}}\left\{\begin{array}[]{ll}(\nu/\overline{\nu})^{p_{1}},&\nu<\overline{\nu}\vspace{0.2cm}\\ (\nu/\overline{\nu})^{p_{2}},&\nu\geq\overline{\nu}\\ \end{array}\right.$ (10)
    Here $\overline{\nu}\sim 10-30\,{\rm GHz}$, $\mathcal{L}_{\rm nb}\sim 10^{38}-10^{39}\,{\rm erg/s}$, $p_{1}\sim 0.6-2.0$ and $p_{2}\sim-0.5$ for the three FRBs currently detected with a confirmed PRS (Chatterjee et al., 2017; Niu et al., 2022; Bruni et al., 2023). The luminosity distance $d_{L}\sim 400-1200\,{\rm Mpc}$ is determined by the known redshift $z\sim 0.1-0.2$.
    We assume that the bolometric luminosity for nebula $\mathcal{L}_{\rm nb}$ peaks in the radio bands, to estimate the minimum required energy ${\rm Min}(\mathcal{E}_{e}^{\rm nb})$ stored in the electrons. The LF for radio-emitting electron is $\gamma_{e}\approx(4\pi m_{e}c\overline{\nu}/3eB_{\rm nb})^{1/2}$ and the energy loss rate is $P_{e}\approx(4/3)\sigma_{T}c\gamma_{e}^{2}B_{\rm nb}^{2}/8\pi$. The total number of electrons in the magnetised nebula is roughly $N_{e}\approx\mathcal{L}_{\rm nb}/P_{e}$ and the minimum required energy in the nebula is then (see Kashiyama & Murase 2017)
@@
  • Constraint on source DM: The number density of free electrons in the SN ejecta and magnetised nebula are given by $n_{\rm e,ej}\approx 3M_{\rm ej}/(4\pi R_{\rm ej}^{3}\mu_{e}\overline{A}m_{\rm H})$ and $n_{\rm e,nb}\approx 3M_{\rm nb}/(4\pi R_{\rm nb}^{3}\mu_{e}\overline{A}m_{\rm H})$, respectively. Here $M_{\rm ej/nb}$ is the ejecta/nebula mass, $\overline{A}=10$ is the mean atomic mass number and $\mu_{e}\approx 1$ for the singly ionized state corresponding to electron temperature $\mathcal{T}_{e}\sim 10^{4}\,{\rm K}$ (see Kashiyama & Murase 2017). The total DM contribution close to the source, from the SN ejecta and magnetised wind nebula is
@@
    where $R_{\rm nb}$ and $R_{\rm ej}$ are obtained by solving equations (6) and (7). Although $M_{\rm nb}$ is uncertain, a reasonable lower limit can be obtained using the Goldreich-Julian (GJ) density (Goldreich & Julian, 1969)
    $M_{\rm nb}/t_{\rm age}\gtrsim\dot{M}_{\rm GJ}\approx(2.4\times 10^{-11}\,M_{\odot}/{\rm s})\,\mu_{\pm,6}B_{\rm int,15}P_{i,-3}^{-2},$ (13)
    where $\mu_{\pm}\sim 10^{5-6}$ is the pair multiplicity defined as the ratio of the pair density to the Goldreich-Julian (GJ) density. The ejecta DM contribution from the SNR should evolve over time as ${\rm DM}_{\rm ej}\propto t^{-2}$. The NS needs to be old enough such that ${\rm DM}_{\rm ns}\approx n_{\rm e,ej}R_{\rm ej}+n_{\rm e,nb}R_{\rm nb}\lesssim{\rm DM}_{\rm host}$.
@@
    is also less than unity, for a energy-dependent SSA cross section $\sigma_{\rm sa}$ (Ghisellini & Svensson, 1991).
  • Size constraint for FRB source: Lastly, the size of the magnetised nebula should not exceed the observed upper limit on the size of the PRS, given by its angular size and luminosity distance for the FRB. From VLBI and VLA observations of FRBs 121102, 190520 and 201124, $R_{\rm nb}\lesssim 1-10\,{\rm pc}$ (see Tendulkar et al. 2017; Niu et al. 2022; Bruni et al. 2023).
@@ 4.2 Implications for FRBs detected with PRS
To compute the synchrotron emission from magnetized nebula, we first need to constrain the NS parameters including $B_{\rm dip}\sim 10^{12}-10^{15}\,{\rm G}$, $P_{i}\sim 1-30\,{\rm ms}$ and $t_{\rm age}\gtrsim t_{\rm obs}$, where $t_{\rm obs}$ is the time since initial detection of the FRB source. For our analysis, we consider two progenitors: (a) an ultra-stripped SNe with $M_{\rm ej}\sim 0.1\,M_{\odot}$ and $\mathcal{E}_{\rm SN}\sim 10^{50}\,{\rm erg}$ (USSN model), and (b) a conventional core-collapse SNe with $M_{\rm ej}\sim 3.0\,M_{\odot}$ and $\mathcal{E}_{\rm SN}\sim 10^{51}\,{\rm erg}$ (CCSN model). For the injected magnetic energy (see equation 2), we adopt $B_{\rm int}=10^{16}\,{\rm G}$ and $t_{\rm inj}=0.6\,{\rm yr}$ as in Margalit & Metzger (2018). We discuss the combination of NS parameters $B_{\rm dip}$, $P_{i}$ and $t_{\rm age}$ for which the magnetar satisfies all the necessary constraints as discussed in Section 4.1.
@@
Figure 2 shows the constraints on $P_{i}-B_{\rm dip}$ parameter space for the nebula energy requirement in red, DM contribution from the ejecta and nebula in green, and the NS spindown luminosity in black. The results for FRB 121102 (R1), FRB 190520 (R2) and FRB 201124 (R3) are shown in the top, middle and bottom row panels, respectively, with solid/dashed curves corresponding to the USSN/CCSN model in each panel. Along with the NS spindown luminosity, we also consider magnetic energy injection with $B_{\rm int}=10^{16}\,{\rm G}$ and $t_{\rm inj}=0.6\,{\rm yr}$ for all cases. We find that the DM criteria for SN ejecta and MWN contribution (${\rm DM}_{\rm ej+nb}<{\rm DM}_{\rm host}$), and the MWN minimum energy requirement ($\varepsilon_{\rm nb,min}<\varepsilon_{\rm inj,max}$) are the most constraining. In the left, center and right column panels, we vary the NS age: $t_{\rm age}\sim 10-40\,{\rm yr}$ for R1 and R2, and $t_{\rm age}\sim 5-20\,{\rm yr}$ for R3. With an increase in $t_{\rm age}$, $DM_{\rm nb+ej}$ decreases faster due to expansion as compared to the reduction in $E_{\rm inj,max}$ from adiabatic energy losses, thereby allowing a larger $P_{i}-B_{\rm dip}$ parameter space. For a given $P_{i}$, ${\rm DM}_{\rm ej+nb}$ reduces faster for a larger $B_{\rm dip}$, as more energy gets injected into the SN ejecta within a shorter timescale leading to a faster ejecta expansion. However, a smaller $B_{\rm dip}$ is allowed for a given $P_{i}$ to prevent considerable adiabatic losses.
@@
For a given $t_{\rm age}$, it is easier to satisfy the DM criterion (${\rm DM}_{\rm ej+nb}<{\rm DM}_{\rm host}$) in the order R2 > R1 > R3, which directly follows from the inferred $DM_{\rm host}$ values for these three localized FRBs. However, the energy criterion ($\varepsilon_{\rm nb,min}<\varepsilon_{\rm inj,max}$) is satisfied in a larger $P_{i}-B_{\rm dip}$ parameter space for R3 as compared to the other two sources, primarily due to smaller $q_{1}$ leading to more energy injection near the peak of the spectrum (see equation 10). Although the qualitative trends in parameter constraints obtained from the near-source DM and nebula energy criteria are effectively same, for both USSN and CCSN progenitors at a given $t_{\rm age}$, the DM (energy) constraint is harder (easier) to satisfy for the CCSN case due to their larger $M_{\rm ej}$ ($E_{\rm sn})$. For $t_{\rm age}\gtrsim 10\,{\rm yr}$ and USSN (CCSN) progenitor, the allowed parameter space satisfying both these conditions is limited to: $P_{i}\sim 1-10\,{\rm ms}$ and $B_{\rm dip}\sim 10^{12}-10^{14}\,{\rm G}$ ($P_{i}\sim 1-2.5\,{\rm ms}$ and $B_{\rm dip}\sim 5\times 10^{12}-2\times 10^{14}\,{\rm G}$) for R1, $P_{i}\sim 1-10\,{\rm ms}$ and $B_{\rm dip}\sim 10^{12}-10^{14}\,{\rm G}$ ($P_{i}\sim 1-20\,{\rm ms}$ and $B_{\rm dip}\sim 10^{12}-2\times 10^{14}\,{\rm G}$) for R2, and $P_{i}\sim 1-20\,{\rm ms}$ and $B_{\rm dip}\sim 10^{12}-10^{14}\,{\rm G}$ ($P_{i}\sim 1-3\,{\rm ms}$ and $B_{\rm dip}\sim 5\times 10^{12}-3\times 10^{14}\,{\rm G}$) for R3. For both USSN and CCSN cases, the parameter space satisfying $L_{\rm sd}\gtrsim 10^{42}\,{\rm erg/s}$ generally overlaps with that allowed by ${\rm DM}_{\rm ej+nb}<{\rm DM}_{\rm host}$ and $\varepsilon_{\rm nb,min}<\varepsilon_{\rm inj,max}$ criteria, especially for $t_{\rm age}\gtrsim 10\,{\rm yr}$. This indicates that the NS spindown luminosity is likely to be the primary energy source for these three FRBs.

### v2

@@ 2.2.1 Energy injection
Both rotational energy injection (Dai et al., 2017; Kashiyama and Murase, 2017) and magnetic energy injection (Beloborodov, 2017; Margalit and Metzger, 2018; Zhao and Wang, 2021) models have been proposed to explain the PRS luminosity as well as the large RM detected for FRB 20121102A (Michilli et al., 2018). The observed persistent radio emission from FRBs can be interpreted as synchrotron emission from electron-positron pairs in the nebula.
@@
${\rm Max}(\mathcal{E}^{\rm inj}_{e})\approx\epsilon_{e}\mathcal{N}_{\rm sd}(t_{\rm age})\int_{0}^{t_{\rm age}}L_{\rm sd}(t^{\prime})dt^{\prime}$
$+\ \epsilon_{e}\mathcal{N}_{\rm mag}(t_{\rm age})\int_{0}^{t_{\rm age}}L_{\rm mag}(t^{\prime})dt^{\prime}.$ (3)
@@ 2.2.2 Particle number distribution
The rotational or magnetic energy is injected into the nebula together with the particles. The injected particles are accelerated to relativistic energies inside or around the termination shock before entering the nebula. The lepton injection rate $\dot{n}_{\mathcal{E}_{e}}^{\rm inj}$ is given by a broken power-law function (see e.g, Tanaka and Takahara 2010; Murase et al. 2015)
[…]
where $R_{\rm nb}$ is the nebula radius, $\mathcal{R}_{0}\sim(2-q_{1})^{-1}+(q_{2}-2)^{-1}\sim 5$ is the correction factor for the lepton normalization, $q_{1}<2$ and $q_{2}\geq 2$ are the low- and high-energy spectral indices. For our analysis, we adopt the characteristic LF of the accelerated leptons to be $\gamma_{b}\sim 10^{5}$ for rotation-powered and $\gamma_{b}\sim 10^{3}$ for magnetic-powered models. The spectral indices are $q_{1}=1.5$ and $q_{2}=2.5$ for rotation-powered scenario, and $q_{1}=q_{2}=2.0$ for magnetic-powered case.
We assume $10^{3}\leq\gamma_{e}\leq 10^{7}$, based on the multi-wavelength modelling of young Galactic PWNe (Tanaka and Takahara, 2013). The effective pair multiplicity can be generally expressed in terms of $\gamma_{b}$, $q_{1}$ and $q_{2}$ (see Murase et al. 2015). The characteristic Lorentz factor of accelerated electrons and positrons is adopted within the typical range inferred from relativistic shock and reconnection simulations. This choice is intended as an empirical representation of the injection process, rather than a direct outcome of detailed magnetohydrodynamic (MHD) evolution. Consequently, our formulation captures the broad energetics of the post-shock nebula without explicitly solving the forward–reverse shock dynamics for arbitrary magnetization.
@@ 3.1 MWN and SNR evolution
The density profile of the SN ejecta can be described using a power-law function $\rho_{\rm ej}=(3-\delta)M_{\rm ej}/4\pi R_{\rm ej}^{3}(R/R_{\rm ej})^{-\delta}$ (Chevalier and Soker, 1989; Kasen and Bildsten, 2010), where $\delta=1$ is the fiducial value of the index, $M_{\rm ej}$ is the ejecta mass and $R_{\rm ej}$ is the ejecta radius. The time evolution of the SN ejecta internal energy is given by
@@ 3.2 Synchrotron emission
The average magnetic field in the nebula is given by $B_{\rm nb}=\sqrt{6\mathcal{E}_{B}/R_{\rm nb}^{3}}$, where $\mathcal{E}_{B}$ is the magnetic energy in the nebula. The magnetic energy in the nebula evolves as (see e.g., Murase et al. 2016, 2021)
@@
In this work, we consider $c_{B}=0$ in the limit when energy loss due to adiabatic expansion is negligible, as done previously for Galactic PWNe (Tanaka and Takahara, 2010) and pulsar-powered PWNe (Murase et al., 2015, 2016).
Here the post-shock energy densities are parameterized using $\epsilon_{e}$ and $\epsilon_{B}$, which describe the downstream energy partition between relativistic particles and magnetic fields, respectively. These quantities are treated phenomenologically rather than derived self-consistently from the upstream wind magnetization $\sigma$. Implicit in our approach is that a fraction $0<\eta_{\rm diss}<1$ of the upstream Poynting flux is efficiently converted into the downstream internal energy of particles. We assume that $\eta_{\rm diss}$ is sufficiently large for the downstream energy fractions ($\epsilon_{B}$, $\epsilon_{e}$) to be physically reasonable, implying that a significant portion of the Poynting flux is locally dissipated. This can be achieved through various mechanisms, including magnetic reconnection in a striped wind (Coroniti, 1990; Cerutti et al., 2020), dissipation at termination shock (Kennel and Coroniti, 1984), current-driven kink instabilities (Mizuno et al., 2011), and turbulence in the nebula (Porth et al., 2014; Zrake and Arons, 2017). As the efficiencies of these processes remain uncertain and model-dependent, a detailed treatment of the dissipation physics is beyond the scope of this work.
We compute the synchrotron emission from the nebula by solving kinetic equations for the photons and electrons and positrons (see e.g., Murase et al. 2021)
@@
where $t_{\rm comp}^{\rm nb}$ is the Comptonisation timescale in the nebula, $t_{\rm esc}^{\rm nb}=R_{\rm nb}/c$ is the photon escape timescale, $t_{\gamma\gamma}$ is the photon annihilation timescale, $\partial n_{\mathcal{E}_{\gamma}}^{\rm IC/syn}/\partial t$ is the photon generation rate from IC/synchrotron process. $P_{\rm ad}$, $P_{\rm syn}$ and $P_{\rm IC}$ are the energy-loss rates due to adiabatic expansion, synchrotron radiation and IC, respectively, $\partial n_{\mathcal{E}_{e}}^{\gamma\gamma}/\partial t$ is the electron-positron production rate from photon annihilation and $\dot{n}_{\mathcal{E}_{e}}^{\rm inj}$ is the electron injection rate from equation (4).
@@ 4.1 Constraints on model parameters
Figure 2: Constraints on NS parameters $B_{\rm dip}$ and $P_{i}$ from the nebula energy requirement (red curves), ejecta and nebula contribution to the source DM (green curves), and the NS spindown luminosity (black curves) are shown for FRB 20121102A/20190520B/20201124A in top/middle/bottom row. In each panel, the solid curves show the results for ($E_{\rm sn}$, $M_{\rm ej}$) = ($10^{50}\,{\rm erg}$, $0.1\,M_{\odot}$) while the dashed curves show results for ($E_{\rm sn}$, $M_{\rm ej}$) = ($10^{51}\,{\rm erg}$, $3.0\,M_{\odot}$). For the magnetic energy injection, we fix $B_{\rm int}=10^{16}\,{\rm G}$ and $t_{\rm inj}=0.6\,{\rm yr}$ for all cases. For FRBs 20121102A and 20190520B, we vary $t_{\rm age}=10\,{\rm yr}$ (left panel), $20\,{\rm yr}$ (middle panel) and $40\,{\rm yr}$ (right panel) as $t_{\rm obs}>5\,{\rm yr}$. In case of FRB 20201124A, results are shown for $t_{\rm age}=5\,{\rm yr}$ (left panel), $10\,{\rm yr}$ (middle panel) and $20\,{\rm yr}$ (right panel).
  • Magnetised nebula energy requirement: The minimum energy in relativistic electrons and positrons required to explain the quasi-steady radio emission from FRB sources should be less than the maximum energy stored in the nebula which is given by ${\rm Max}(\mathcal{E}_{e}^{\rm inj})$ (see e.g., equation 2.2.1). The observed quasi-steady radio spectrum for FRBs can be fitted by
[…]
    Here $\overline{\nu}\sim 10-30\,{\rm GHz}$, $\mathcal{L}_{\rm nb}\sim 10^{38}-10^{39}\,{\rm erg/s}$, $p_{1}\sim 0.6-2.0$ and $p_{2}\sim-0.5$ for the five FRBs currently detected with a confirmed PRS (Chatterjee et al., 2017; Niu et al., 2022; Bruni et al., 2023, 2024; Moroianu et al., 2025). The luminosity distance $d_{L}\sim 400-1200\,{\rm Mpc}$ is determined by the known redshift $z\sim 0.1-0.2$.
    We assume that the bolometric luminosity for nebula $\mathcal{L}_{\rm nb}$ peaks in the radio bands, to estimate the minimum required energy ${\rm Min}(\mathcal{E}_{e}^{\rm nb})$ stored in the electrons and positrons. The LF for radio-emitting electron is $\gamma_{e}\approx(4\pi m_{e}c\overline{\nu}/3eB_{\rm nb})^{1/2}$ and the energy loss rate is $P_{e}\approx(4/3)\sigma_{T}c\gamma_{e}^{2}B_{\rm nb}^{2}/8\pi$. The total number of electrons and positrons in the magnetised nebula is roughly $N_{e}\approx\mathcal{L}_{\rm nb}/P_{e}$ and the minimum required energy in the nebula is then (see Kashiyama and Murase 2017)
@@
  • Constraint on source DM: The number density of free electrons in the SN ejecta and magnetised nebula are given by $n_{\rm e,ej}\approx 3M_{\rm ej}/(4\pi R_{\rm ej}^{3}\mu_{e}\overline{A}m_{\rm H})$ and $n_{\rm e,nb}\approx 3M_{\rm nb}/(4\pi R_{\rm nb}^{3}\mu_{e}\overline{A}m_{\rm H})$, respectively. Here $M_{\rm ej/nb}$ is the ejecta/nebula mass, $\overline{A}=10$ is the mean atomic mass number and $\mu_{e}\approx 1$ for the singly ionized state corresponding to electron temperature $\mathcal{T}_{e}\sim 10^{4}\,{\rm K}$ (see Kashiyama and Murase 2017). The total DM contribution close to the source, from the SN ejecta and magnetised wind nebula is
@@
    where $R_{\rm nb}$ and $R_{\rm ej}$ are obtained by solving equations (6) and (7). Although $M_{\rm nb}\gtrsim 10^{-9}-10^{-5}\,M_{\odot}$ is largely uncertain (Murase et al., 2015), a reasonable lower limit can be obtained using the Goldreich-Julian (GJ) density (Goldreich and Julian, 1969)
    $M_{\rm nb}/t_{\rm age}\gtrsim\dot{M}_{\rm\pm}\approx(2.4\times 10^{-11}\,M_{\odot}/{\rm s})\,\mu_{\pm,6}B_{\rm int,15}P_{i,-3}^{-2},$ (13)
    where $\dot{M}_{\rm\pm}$ is the mass-loss rate due to electron-positron pairs and $\mu_{\pm}\sim 10^{5-6}$ is the pair multiplicity defined as the ratio of the pair density to the Goldreich-Julian (GJ) density. It is difficult to infer a plausible theoretical value for the pair multiplicity from first principles. In this study, we adopt $M_{\rm nb}\sim 10^{-7}\,M_{\odot}$ as a fiducial value (Murase et al., 2015). The ejecta DM contribution from the SNR should evolve over time as ${\rm DM}_{\rm ej}\propto t^{-2}$. The NS needs to be old enough such that ${\rm DM}_{\rm ns}\approx n_{\rm e,ej}R_{\rm ej}+n_{\rm e,nb}R_{\rm nb}\lesssim{\rm DM}_{\rm host}$.
@@
    is also less than unity, for a energy-dependent SSA cross section $\sigma_{\rm sa}$ (Ghisellini and Svensson, 1991).
  • Size constraint for FRB source: Lastly, the size of the magnetised nebula should not exceed the observed upper limit on the size of the PRS, given by its angular size and luminosity distance for the FRB. From VLBI and VLA observations of FRBs 20121102A, 20190520B and 20201124A, $R_{\rm nb}\lesssim 1-10\,{\rm pc}$ (see Tendulkar et al. 2017; Niu et al. 2022; Bruni et al. 2023).
@@ 4.2 Implications for FRBs detected with PRS
To compute the synchrotron emission from magnetized nebula, we first need to constrain the NS parameters including $B_{\rm dip}\sim 10^{12}-10^{15}\,{\rm G}$, $P_{i}\sim 1-30\,{\rm ms}$ and $t_{\rm age}\gtrsim t_{\rm obs}$, where $t_{\rm obs}$ is the time since initial detection of the FRB source. For our analysis, we consider two progenitors: (a) an ultra-stripped SNe with $M_{\rm ej}\sim 0.1\,M_{\odot}$ and $\mathcal{E}_{\rm SN}\sim 10^{50}\,{\rm erg}$ (USSN model), and (b) a conventional core-collapse SNe with $M_{\rm ej}\sim 3.0\,M_{\odot}$ and $\mathcal{E}_{\rm SN}\sim 10^{51}\,{\rm erg}$ (CCSN model). For the injected magnetic energy (see equation 2), we adopt $B_{\rm int}=10^{16}\,{\rm G}$ and $t_{\rm inj}=0.6\,{\rm yr}$ as in Margalit and Metzger (2018). We discuss the combination of NS parameters $B_{\rm dip}$, $P_{i}$ and $t_{\rm age}$ for which the magnetar satisfies all the necessary constraints as discussed in Section 4.1.
@@
Figure 2 shows the constraints on $P_{i}-B_{\rm dip}$ parameter space for the nebula energy requirement in red, DM contribution from the ejecta and nebula in green, and the NS spindown luminosity in black. The results for FRBs 20121102A, 20190520B and 20201124A are shown in top, middle and bottom row panels, respectively, with solid/dashed curves corresponding to USSN/CCSN model in each panel. Along with NS spindown luminosity, we also consider magnetic energy injection with $B_{\rm int}=10^{16}\,{\rm G}$ and $t_{\rm inj}=0.6\,{\rm yr}$ for all cases. We find that the DM criteria for SN ejecta and MWN contribution (${\rm DM}_{\rm ej+nb}<{\rm DM}_{\rm host}$), and the MWN minimum energy requirement ($\varepsilon_{\rm nb,min}<\varepsilon_{\rm inj,max}$) are the most constraining. In the left, center and right column panels, we vary the NS age: $t_{\rm age}\sim 10-40\,{\rm yr}$ for FRBs 20121102A and 20190520B, and $t_{\rm age}\sim 5-20\,{\rm yr}$ for FRB 20201124A. With an increase in $t_{\rm age}$, $DM_{\rm nb+ej}$ decreases faster due to expansion as compared to the reduction in $E_{\rm inj,max}$ from adiabatic energy losses, thereby allowing a larger $P_{i}-B_{\rm dip}$ parameter space. For a given $P_{i}$, ${\rm DM}_{\rm ej+nb}$ reduces faster for a larger $B_{\rm dip}$, as more energy gets injected into the SN ejecta within a shorter timescale leading to a faster ejecta expansion. However, a smaller value of $B_{\rm dip}$ is allowed for a given $P_{i}$ to prevent considerable adiabatic losses.
@@
For a given $t_{\rm age}$, it is easier to satisfy the DM criterion (${\rm DM}_{\rm ej+nb}<{\rm DM}_{\rm host}$) in the order FRB 20190520B > FRB 20121102A > FRB 20201124A, which directly follows from the inferred $DM_{\rm host}$ values for these three localized sources. However, the energy criterion ($\varepsilon_{\rm nb,min}<\varepsilon_{\rm inj,max}$) is satisfied in a larger $P_{i}-B_{\rm dip}$ parameter space for FRB 20201124A as compared to the other two sources, primarily due to smaller $q_{1}$ leading to more energy injection near the peak of the spectrum (see equation 10). Although the qualitative trends in parameter constraints obtained from near-source DM and nebula energy criteria are effectively same, for both USSN and CCSN progenitors at a given $t_{\rm age}$, the DM (energy) constraint is harder (easier) to satisfy for the CCSN case due to their larger $M_{\rm ej}$ ($E_{\rm sn})$. For $t_{\rm age}\gtrsim 10\,{\rm yr}$ and USSN (CCSN) progenitor, the allowed parameter space satisfying both these conditions is limited to: $P_{i}\sim 1-10\,{\rm ms}$ and $B_{\rm dip}\sim 10^{12}-10^{14}\,{\rm G}$ ($P_{i}\sim 1-2.5\,{\rm ms}$ and $B_{\rm dip}\sim 5\times 10^{12}-2\times 10^{14}\,{\rm G}$) for FRB 20121102A, $P_{i}\sim 1-10\,{\rm ms}$ and $B_{\rm dip}\sim 10^{12}-10^{14}\,{\rm G}$ ($P_{i}\sim 1-20\,{\rm ms}$ and $B_{\rm dip}\sim 10^{12}-2\times 10^{14}\,{\rm G}$) for FRB 20190520B, and $P_{i}\sim 1-20\,{\rm ms}$ and $B_{\rm dip}\sim 10^{12}-10^{14}\,{\rm G}$ ($P_{i}\sim 1-3\,{\rm ms}$ and $B_{\rm dip}\sim 5\times 10^{12}-3\times 10^{14}\,{\rm G}$) for FRB 20201124A. For both USSN and CCSN cases, the parameter space satisfying $L_{\rm sd}\gtrsim 10^{42}\,{\rm erg/s}$ overlaps with that allowed by ${\rm DM}_{\rm ej+nb}<{\rm DM}_{\rm host}$ and $\varepsilon_{\rm nb,min}<\varepsilon_{\rm inj,max}$, especially for $t_{\rm age}\gtrsim 10\,{\rm yr}$. This indicates that NS spindown luminosity is likely to be the primary energy source for these three FRBs.